Solve the equation. Round the solutions to the nearest hundredth, if necessary.
-t^2 + 2t + 15 = 0

Found 2 solutions by LinnW, richwmiller:
Answer by LinnW(1048)

-t^2 + 2t + 15 = 0
multiply each side by -1
t^2 -2t -15 = 0
(t - 5)(t + 3) = 0
So t = 5 or t = -3

Answer by richwmiller(17219)

multiply by -1
t^2-2t-15=0
find factors of -15 which add up to -2
-5 and 3
(t-5)*(t+3)=0
t=5
t=-3